WHAT IS THE MOLALITY (m) OF A 250 g OF SOLUTION THAT CONTAINS 7.3 GRAMS OF GLUCOSE C6H12O6? WHAT IS THE MOLECULAR WEIGHT OF SOLUTE?
1
Expert's answer
2020-11-10T14:05:46-0500

M(C6H12O6)=180.16 g/mol

n(C6H12O6)=m/M = 7.3/180.16=0.04 mol

Cm(C6H12O6)=0.04/0.25=0.16 m